# Sliced Rényi Pufferfish Privacy: Directional Additive Noise Mechanism and Private Learning with Gradient Clipping

### v2

@@ Abstract
We study privatization mechanism design and privacy accounting in the Pufferfish family, addressing two practical gaps of Rényi Pufferfish Privacy (RPP): high-dimensional optimal transport (OT) calibration and the absence of a general, mechanism-agnostic composition rule for iterative learning. We introduce Sliced Rényi Pufferfish Privacy (SRPP), which replaces high-dimensional comparisons by directional ones over a set of unit vectors, enabling geometry-aware and tractable guarantees. To calibrate noise without high-dimensional OT, we propose sliced Wasserstein mechanisms that compute per-direction (1-D) sensitivities, yielding closed-form, statistically stable, and anisotropic calibrations. We further define SRPP Envelope (SRPE) as computable upper bounds that are tightly implementable by these sliced Wasserstein mechanisms. For iterative deep learning algorithms, we develop a decompose-then-compose SRPP-SGD scheme with gradient clipping based on a History-Uniform Cap (HUC), a pathwise bound on one-step directional changes that is uniform over optimization history, and a mean-square variant (ms-HUC) that leverages subsampling randomness to obtain on-average SRPP guarantees with improved utility. The resulting HUC and ms-HUC accountants aggregate per-iteration, per-direction Rényi costs and integrate naturally with moments-accountant style analyses. Finally, when multiple mechanisms are trained and privatized independently under a common slicing geometry, our analysis yields graceful additive composition in both worst-case and mean-square regimes. Our experiments indicate that the proposed SRPP-based methods achieve favorable privacy–utility trade-offs in both static and iterative settings.
@@ 1 Introduction
Differential Privacy (DP) dwork2006calibrating is the de facto standard for measuring data privacy risk. It treats each individual record as a secret and requires that the output distribution of a randomized algorithm change only slightly when a single record is added or removed. DP and its variants, such as Rényi DP (RDP) Mironov2017; papernot2021hyperparameter and zCDP bun2016concentrated, enjoy graceful composition properties that quantify cumulative privacy loss when the same (or related) datasets are used multiple times. These properties underpin DP-SGD Abadi2016, a canonical decompose-then-compose approach to private deep learning: each iteration applies a subsampled Gaussian mechanism with per-example gradient clipping and Gaussian noise, and a moments accountant aggregates per-iteration RDP into an overall $(\epsilon,\delta)$ guarantee.
By adopting a similar worst-case probabilistic indistinguishability, Pufferfish Privacy (PP) kifer2014pufferfish generalizes DP by allowing secrets beyond individual records and by modeling probabilistic relationships between secrets and datasets via priors on an adversary’s beliefs. Song et al. Song2017 provide the first general-purpose privatization mechanism for PP, the Wasserstein mechanism, which uses the $\infty$-Wasserstein distance as an analogue of sensitivity and generalizes the Laplace mechanism. Ding ding2022kantorovich proposes a Kantorovich-based exponential (and Gaussian) additive-noise mechanism for PP calibrated to 1-Wasserstein sensitivity. However, computing $\infty$-Wasserstein distances is statistically and computationally challenging in high dimensions, and tractable, general design principles for PP mechanisms remain largely open jajodia2025encyclopedia.
A second central limitation of PP is the lack of practical, general-purpose composition theorems for privacy accounting and iterative algorithms like SGD. Existing results require specific structural assumptions: universally composable evolution scenarios kifer2014pufferfish assume deterministic secret-dataset relationships, while the Markov Quilt Mechanism Song2017 applies only to Bayesian networks. Recent work zhang2024differential explains this limitation fundamentally: the probabilistic secret-dataset relationship induces dependencies among mechanisms, breaking standard composition. Their proposed inverse composition requires solving an infinite-dimensional convex optimization problem, which is is not straightforward to apply within iterative learning algorithms such as SGD.
@@
Rényi Pufferfish Privacy (RPP) pierquin2024renyi partially addresses these issues by recasting PP in terms of Rényi divergences and introducing the General Wasserstein Mechanism (GWM), which uses the shift-reduction lemma feldman2018privacy; altschuler2022privacy to calibrate additive noise via $\infty$-Wasserstein distance. Extensions include an approximate version (GAWM) for improved utility and a distribution-aware variant (DAGWM) using $p$-Wasserstein distances to capture average-case data geometry. While these Wasserstein-based mechanisms significantly advance the PP framework, they inherit the curse of dimensionality from optimal transport: estimating Wasserstein distances and sensitivities remains statistically and computationally expensive in high dimensions, limiting practical applicability.
To address composition, RPP proves Privacy Amplification by Iteration (PABI) results for contractive noisy processes pierquin2024renyi, enabling the first use of PP-style guarantees in private convex optimization. However, PABI requires contractivity of the update map, restricting applicability to non-convex objectives and non-contractive optimizers like momentum-based methods. Moreover, PABI still depends on Wasserstein distances between one-step updates, reintroducing high-dimensional optimal transport challenges. Consequently, PABI does not provide a universal, mechanism-agnostic composition theory for Pufferfish.
In this paper, we address these limitations by taking a directional, sliced view of RPP and by designing mechanisms and accountants that avoid high-dimensional optimal transport and contractivity assumptions. At a high level, we (i) introduce sliced RPP notions and sliced Wasserstein mechanisms that reduce high-dimensional calibration to one-dimensional problems, (ii) develop HUC- and ms-HUC–based SRPP-SGD schemes that support practical privacy accounting for subsampled gradient methods, and (iii) establish additive composition theorems for collections of SRPP/ms-SRPP mechanisms sharing a common slicing geometry. Our contributions are summarized as follows.
  • We introduce two notions of Sliced Rényi Pufferfish Privacy (SRPP), average SRPP (Ave-SRPP) and joint SRPP (Joint-SRPP), as directional refinements of RPP that compare secrets along a fixed set of unit vectors. Ave-SRPP controls the mean Rényi distinguishability across slices, while Joint-SRPP controls the Rényi distinguishability of the joint pair of slice index and projection, penalizing rare high-risk directions that could be hidden by averaging. Having both notions lets us trade off utility and robustness to directional outliers while keeping the geometry-aware, tractable structure of SRPP.
  • We propose sliced Wasserstein mechanisms, which replace $d$-dimensional Wasserstein sensitivities with per-direction (1D) sensitivities, and define SRPP Envelope (SRPE) as computable upper bounds that are tightly implementable by these mechanisms, yielding explicit privacy certificates for Ave-/Joint-SRPP.
  • For iterative learning with gradient clipping, we develop SRPP-SGD based on a History-Uniform Cap (HUC), which provides additive per-iteration, per-direction Rényi bounds and closed-form Gaussian calibration, and a mean-square variant (ms-HUC) together with ms-SRPP that leverage subsampling randomness to obtain on-average guarantees with improved utility. The resulting HUC and ms-HUC accountants integrate naturally with moments-accountant–style analyses and avoid both contraction assumptions and high-dimensional Wasserstein estimates, accommodating non-contractive optimizers such as SGD and Adam.
  • When multiple mechanisms are trained independently and privatized via our SRPP/ms-SRPP schemes under a common slicing geometry, we prove graceful additive composition: the total Ave-/Joint- or ms-Ave/ms-Joint-SRPP cost is bounded by the sum of the per-mechanism costs, enabling modular, plug-and-play privacy accounting for pipelines and cascades.

### v3

# Sliced Rényi Pufferfish Privacy: Tractable Privatization Mechanism and Private Learning with Gradient Clipping

@@ Abstract
We study the design of a privatization mechanism and privacy accounting in the Pufferfish Privacy (PP) family. Specifically, motivated by the curse of dimensionality and lack of practical composition tools for iterative learning in the recent Rényi Pufferfish Privacy (RPP) framework, we propose Sliced Rényi Pufferfish Privacy (SRPP). SRPP preserves PP/RPP semantics (customizable secrets with probability-aware secret–dataset relationships) while replacing high-dimensional Rényi divergence with projection-based quantification via two sliced measures, Average SRPP and Joint SRPP. We develop sliced Wasserstein mechanisms, yielding sound SRPP certificates and closed-form Gaussian noise calibration. For iterative learning systems, we introduce an SRPP-SGD scheme with gradient clipping and new accountants based on History-Uniform Caps (HUC) and a subsampling-aware variant (sa-HUC), enabling decompose-then-compose privatization and additive composition under a common slicing geometry. Experiments on static and iterative privatization show that the proposed framework exhibits favorable privacy–utility trade-offs, as well as practical scalability.
@@ 1 Introduction
Differential Privacy (DP) dwork2006calibrating is the de facto standard for measuring data privacy risk. Under DP semantics, the secret is an individual record’s presence or absence, and an adversary with arbitrary side information cannot draw substantially different conclusions from the released output about that inclusion. DP quantifies risk via $\epsilon$-DP or $(\epsilon,\delta)$-DP, i.e., worst-case indistinguishability between neighboring datasets, equivalently expressed through bounds on max divergence and hockey-stick divergence, respectively. Rényi DP (RDP) Mironov2017; papernot2021hyperparameter preserves DP semantics but measures privacy using $(\alpha,\epsilon)$-Rényi divergence.
Pufferfish Privacy (PP) kifer2014pufferfish generalizes DP semantics by allowing customizable secrets and modeling probabilistic secret–dataset relationships through priors that may represent subjective adversarial beliefs or an objective secret-data-generation process zhang2025differential. $\epsilon$-PP and $(\epsilon,\delta)$-PP Zhang2022 adopt the same indistinguishability measures as DP, while Rényi Pufferfish Privacy (RPP) pierquin2024renyi shares PP semantics but uses $(\alpha,\epsilon)$-Rényi divergence.
Beyond semantics and a quantification measure, a useful privacy framework should provide implementable privatization mechanisms. For DP, Laplace dwork2006calibrating and Gaussian dwork2006our mechanisms calibrate noise to query sensitivity. However, computing sensitivity is generally NP-hard xiao2008output. DP and RDP composition properties underpin DP-SGD with gradient clipping Abadi2016, a decompose-then-compose approach for deep learning without sensitivity computation: each iteration enforces bounded sensitivity via gradient clipping, adds subsampled Gaussian noise, and a moments accountant converts per-iteration RDP costs into an overall $(\epsilon,\delta)$-DP guarantee.
The key technical difference between DP and PP is that PP involves a probabilistic secret-dataset relationship. Because a query only takes the realized dataset as input, PP analysis must account for the induced mapping from secret to output through the prior distributions. Consequently, practical PP calibration depends on prior-driven distributional quantities and is typically carried out from finite samples and estimators, so error-aware, assumption-dependent guarantees are often unavoidable for useful privacy-utility trade-offs.
@@
For Pufferfish privatization, Song et al. Song2017 provide the first general-purpose privatization mechanism, the Wasserstein mechanism, which uses the $\infty$-Wasserstein distance as an analogue of sensitivity and generalizes the Laplace mechanism. Ding ding2022kantorovich proposes a Kantorovich-based exponential (and Gaussian) additive-noise mechanism for PP calibrated to 1-Wasserstein sensitivity. RPP pierquin2024renyi introduces the General Wasserstein Mechanism (GWM), which uses the shift-reduction lemma feldman2018privacy; altschuler2022privacy to calibrate additive noise via $\infty$-Wasserstein distance. Extensions of GWM include a distribution-aware variant (DAGWM) using $p$-Wasserstein distances to capture average-case data geometry. While these Wasserstein mechanisms significantly advance the applicability of the PP family frameworks, they inherit the curse of dimensionality from optimal transport (OT): estimating Wasserstein distances and sensitivities remains statistically and computationally expensive in high dimensions, limiting practical applicability. The tractable general mechanism design principles for PP family remain largely open jajodia2025encyclopedia.
Moreover, PP/RPP lack practical, general-purpose composition theorems for privacy accounting in stochastic optimization algorithms such as SGD. Existing results rely on restrictive structure (e.g., universally composable evolution scenarios kifer2014pufferfish or Bayesian networks via Markov quilts Song2017). Recent work zhang2025differential shows the obstacle is fundamental: probabilistic secret–dataset relationships induce inter-mechanism dependencies that break standard composition; their inverse composition requires solving an infinite-dimensional convex program, making integration with iterative learning nontrivial. RPP’s Privacy Amplification by Iteration (PABI) pierquin2024renyi enables privacy guarantees for iterative convex optimization under contractivity, but still depends on Wasserstein distance computation, reintroducing high-dimensional OT costs.
Therefore, making PP-family guarantees operational in practice requires resolving two fundamental obstacles due to the probabilistic secret-dataset relationship captured by priors: (i) curse of dimensionality in mechanism design, i.e., tractable calibration without high-dimensional Wasserstein computations; and (ii) lack of graceful composition and accounting for iterative optimization and learning, i.e., enabling decompose-then-compose privatization for stochastic iterative methods (e.g., SGD) under PP-family semantics.
Contributions
To address these two obstacles, we propose a new class of privacy definition termed Sliced R’enyi Pufferfish Privacy (SRPP) (Sec.4) together with two complementary, practical privatization mechanisms equipped with rigorous, provable guarantees:
  1. (i)
    a direct privatization mechanism, termed the Sliced Wasserstein Mechanism (SWM) (Sec.4.3), calibrated by an efficiently computable sliced Wasserstein (distance-based) sensitivity (SW-sensitivity) (Sec. 3), which circumvents the curse of dimensionality of OT;
  2. (ii)
    a decompose-then-compose privatization scheme for iterative learning (e.g., SGD) based on History-Uniform Caps (HUC) and their subsampling-aware variant (sa-HUC) under gradient clipping, together with matching subsampling-aware (sa-) SRPP accountants. This avoids Wasserstein sensitivity computation altogether.
For (i), the key idea is to leverage slicing to overcome the scalability bottleneck of OT that arises when calibrating Wasserstein mechanisms via full-dimensional Wasserstein sensitivity (as in RPP’s GWM). Building on the sliced Wasserstein distance rabin2011wasserstein; deshpande2018generative; kolouri2019generalized; nietert2022statistical; bonet2025sliced, we introduce the SW-sensitivity and design the SWM (Sec.4.3), which replaces high-dimensional OT sensitivity with a computable sensitivity surrogate based on one-dimensional (1-D) projections.
To align the privacy quantification with SW-sensitivity, we introduce two SRPP notions, Average SRPP (Sec.4.1) and Joint SRPP (Sec.4.2), that share the same Pufferfish semantics as PP/RPP, but quantify privacy using Average Sliced R’enyi Divergence (Ave-SRD) and Joint Sliced R’enyi Divergence (Joint-SRD), respectively.
For (ii), we develop a general SRPP privatization route for iterative learning via HUC and sa-HUC, enabling DP-SGD-style privatization with gradient clipping without requiring contractivity of the update map. A central element of the contribution is that HUC/sa-HUC isolate a single scenario-dependent quantity—the discrepancy cap (in deterministic form $K_{t}$ and subsampling-aware, mean-square form $K_{t}^{2}$)—that governs the per-iteration secret-induced shift. This identification (i) avoids the utility-draining worst-case "group-privacy" Song2017; pierquin2024renyi bound (see App. A.4 for a formal discussion), and (ii) enables a decompose-then-compose accountant for SGD with gradient clipping, without computing (sliced) Wasserstein sensitivities. Moreover, when multiple mechanisms are trained independently and privatized under a common slicing geometry, we prove graceful additive composition: the total Ave-/Joint- (sa-Ave/sa-Joint-) SRPP cost is bounded by the sum of the per-mechanism costs, yielding modular, plug-and-play privacy accounting for pipelines and cascades.
Experiments on both static query release and iterative learning (SRPP-SGD) demonstrate favorable privacy–utility trade-offs and practical scalability. Formal proofs, detailed discussion and analysis are provided in the appendix.